Exercise 2: The half-life for the first order decomposition of A is 200.0 sec. How much time must elapse for the concentration of A to decrease to (a) one-half; (b) one-sixteenth; (e) one-ninth of its initial concentration?
Answer #1

For a first-order reaction,
t1/2 = 0.693/ K
K = rate constant

we know t1/2 = 200 sec
therefore; K = 0.693/t1/2 = 0.693/200 = 3.465 x 10-3 sec-1

(a) one-half;
t50% left = t1/2 = 200 sec
this is half life time = 200 secs

(b) one-sixteenth (1/16)
t6.25% left = t1/16 = 4 x t1/2 = 4 x 200 sec = 800 sec

(c) one-ninth (1/9)

[A] = [A]0 x e-kt
ln[A]/[A]0 = -KT

therefore;

ln(1/9) = - K x t1/9
t1/9 left = - ln(1/9)/ 3.465 x 10-3 sec-1 = 634 sec